OpenStudy (anonymous):

Simplify -2√7x² • 1/3 √28x The 28x is cubed

OpenStudy (rajee_sam):

\[-2\sqrt{7} x ^{2} \times \frac{ 1 }{ 3 } \sqrt{28} x\] Is this your problem??

OpenStudy (rajee_sam):

x³ at the last

OpenStudy (anonymous):

Yes

OpenStudy (rajee_sam):

\[-2\sqrt{7} . x ^{2} . \frac{ 1 }{3 } . \sqrt{28} . x ^{3}\] Now combine like terms\[-2\sqrt{7} . \sqrt{28} . \frac{ 1 }{3 } . x ^{2} . x ^{3}\] \[-2\sqrt{7} . 2\sqrt{7} . \frac{ 1 }{ 3 } . x ^{2 + 3}\]\[-4 . 7 . \frac{ 1 }{ 3 } . x ^{2 + 3} = -\frac{ 28 }{ 3 } x ^{5}\]

OpenStudy (anonymous):

Thank you 🙌